OpenStudy (anonymous):

The ratio of the perimeters of two similar triangles is 2:3. What is the area of each of these triangles if the sum of their area is 260cm^?

OpenStudy (anonymous):

@Hero @Callisto

OpenStudy (callisto):

Hint:If two figures are similar, (ratio of the sides/perimeter)^2 = ratio of the areas That is \[(\frac{P_1}{P_2})^2 = \frac{A_1}{A_2}\] Now, let \(A_1\) = A, you can express \(A_2\) in terms of A. You can write the equation \(A_1+A_2 = 260\) <- Write \(A_1\) and \(A_2\) in terms of A and solve A.

OpenStudy (anonymous):

@Callisto